A wall of perimeter 38 m surrounds a rectangular yard of length 12 m. What is the width of the yard?
Answer by checkley79(3341)

P=2L+2W
38=2*12+2W
38=24+2W
2W=38-24
2W=14
W=14/2
W=7 M. ANS. FOR THE WIDTH
PROOF:
38=2*12+2*7
38=24+14
38=38